# Non-isotropic translation and modulation invariant Hilbert spaces

### v1

@@ Abstract.
Let $\mathcal{H}$ be a Hilbert space of distributions on $\mathbf{R}^{d}$ which contains at least one non-zero element in $\mathscr{D}^{\prime}(\mathbf{R}^{d})$. If there is a constant $C_{0}>0$ such that
@@ 0. Introduction
$\big{(}\,y\mapsto f(y)\,\big{)}$ $\mapsto\big{(}\,y\mapsto e^{i\langle y,\xi\rangle}f(y)\,\big{)}$
@@
An alternative approach, using the Bargmann transform, to reach similar properties is given by Bais, Pinlodi and Venku Naidu in [1]. In fact, by using their result [1, Theorem 3.1] in combination with some well-known arguments in the distribution theory, one obtains the following improvement of Theorem 0.1.
@@
for some constant $C>0$ in the hypothesis in [1, Theorem 3.1] is absent in Theorem 0.2. Hence Theorem 0.2 is slightly more general than [1, Theorem 3.1].
@@
for some constant $C\geq 1$ which is independent of $f\in\mathcal{B}$ and $x\in\mathbf{R}^{d}$. Then it follows by some standard arguments that $\mathcal{B}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ (see e. g. [15, Proposition 1.5]).
@@ Acknowledgement
The authors are grateful to HRI and the infosys foundations for support for the first author during his visit at HRI in December 2023, making this collaboration possible. The first author is also grateful for the excellent hospitality during his visit at HRI in December 2023. The first author was supported by Vetenskapsrådet (Swedish Science Council) within the project 2019-04890.
@@ 1. Proof of the main result
In this section we prove our main result Theorem 0.3. First we show how the condition (0.2) in [1, Theorem 3.1] can be removed, which leads to Theorem 0.2. Then we prove a lemma, which in combination with Theorem 0.2, essentially leads to Theorem 0.3.
@@
Suppose that $\mathcal{H}$ satisfies the hypothesis in Theorem 0.3. Then there is a constant $C>0$ such that (0.2) holds.
Lemma 1.1 follows by straight-forward continuity arguments. In order to assist the reader here we present the arguments.
Proof.
Let $H_{d}=|x|^{2}-\Delta_{d}$ be the $d$-dimensional harmonic oscillator. Then $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ is the inductive limit of the increasing Hilbert spaces
$\mathcal{H}_{j}\equiv H_{d}^{j}(L^{2}(\mathbf{R}^{d})),\qquad j\in\mathbf{N},$
with norms and scalar products given by
$\|f\|_{\mathcal{H}_{j}}\equiv\|H_{d}^{-j}f\|_{L^{2}(\mathbf{R}^{d})}\quad\text{and}\quad(f,g)_{\mathcal{H}_{j}}\equiv(H_{d}^{-j}f,H_{d}^{-j}g)_{L^{2}(\mathbf{R}^{d})},\quad f,g\in\mathcal{H}_{j}.$
(See e. g. [3, 13].) The fact that $\mathcal{H}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, implies that $\mathcal{H}$ is continuously embedded in $\mathcal{H}_{j}$, provided $j$ is chosen sufficiently large. For such $j$ we have
$|(f,\phi)|=|(f,\phi)_{L^{2}}|\lesssim\|f\|_{\mathcal{H}_{j}}\|H_{d}^{j}\phi\|_{L^{2}}\leq C\|f\|_{\mathcal{H}},$
for some constant $C>0$, and the result follows. ∎
Theorem 0.2 now follows from [1, Theorem 3.1] and Lemma 1.1.
@@
Here the constant $C$ is the same as above. Furthermore, $\|\,\cdot\,\|_{[R]}$ is a norm which arises from the scalar product
@@
exists and is independent of the chosen particular sequence $\{f_{0,j}\}_{j=1}^{\infty}$. Since (1.2) holds for any $f_{0,j}\in\mho$, it follows from the recent estimates and limit properties (1.2) extends to any $f\in\mathcal{H}$. This gives (1).
@@
$\Delta_{R}=\big{(}-x_{j}e_{j}+Q_{R})\setminus Q_{R}\big{)}\bigcup\big{(}Q_{R}\setminus-x_{j}e_{j}+Q_{R}\big{)},$
@@
2. Extensions and variations
We recall that the Pilipović space $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ consists of all Hermite function expansions
@@ 2. Extensions and variations
for some $h>0$ (see [13]). The distribution space $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ can be identified with the set of all formal expansions in (2.1) such that
@@
for every $h>0$. The distribution action between $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ and $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ is then given by
@@
We have now the following extension of Theorem 0.3.
Theorem 2.1.
The proof of Theorem 2.1 follows by a combination of [15, Proposition 1.5], Theorem 0.3, and Proposition 2.2 below. First we need some preparations.
and
@@
with dense and continuous embeddings.
@@
are homeomorphisms. Finally we recall that in terms of the Weyl map

### v2

# Non-isometric translation and modulation invariant Hilbert spaces

@@ Abstract.
Let $\mathcal{H}$ be a Hilbert space, continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, and which contains at least one non-zero element in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$. If there is a constant $C_{0}>0$ such that
@@ 0. Introduction
$f$ $\mapsto e^{i\langle\,\cdot\,,\xi\rangle}f$
@@
An alternative approach, using the Bargmann transform, to reach similar properties is given by Bais, Pinlodi and Venku Naidu in [2]. In fact, by using their result [2, Theorem 3.1] in combination with some well-known arguments in the distribution theory, one obtains the following improvement of Theorem 0.1.
@@
for some constant $C>0$ in the hypothesis in [2, Theorem 3.1] is absent in Theorem 0.2. Hence Theorem 0.2 is slightly more general than [2, Theorem 3.1].
@@
for some constant $C\geq 1$ which is independent of $f\in\mathcal{B}$ and $x\in\mathbf{R}^{d}$. Then it follows by some standard arguments that $\mathcal{B}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ (see e. g. [18, Proposition 1.5]).
@@
The key step for the proof of Theorem 0.3 is to find an equivalent Hilbert norm to $\|\,\cdot\,\|_{\mathcal{H}}$, which is norm preserved under translations and modulations. The result then follows from Theorem 0.2. As a first idea one may try to use the equivalent norm
$\|f\|_{\mathcal{B}}\equiv\sup_{x,\xi\in\mathbf{R}^{d}}\left(\|f(\,\cdot\,-x)e^{i\langle\,\cdot\,,\xi\rangle}\|_{\mathcal{H}}\right).$
A straight-forward control shows that this norm is invariant under translations and modulations. On the other hand, it seems $\|f\|_{\mathcal{B}}$ might fail to be a Hilbert norm, and thereby not being suitable for applying Theorem 0.2.
In our approach to find the sought Hilbert norm, we use some ideas in the construction of $\|\,\cdot\,\|_{\mathcal{B}}$ above, but replace the supremum with mean-values of the form
$\|f\|_{[R]}^{2}\equiv(2R)^{-2d}\iint_{\mathbf{R}^{2d}}\|f(\,\cdot\,-x)e^{i\langle\,\cdot\,,\xi\rangle}\|_{\mathcal{H}}^{2}\,dxd\xi,$
when $R>0$. It follows that each $\|\,\cdot\,\|_{[R]}$ is a Hilbert norm which is uniformly equivalent with $\|\,\cdot\,\|_{\mathcal{H}}$. On the other hand, none of $\|\,\cdot\,\|_{[R]}$ need to be translation nor modulation invariant. However, by increasing $R$, it follows that $\|\,\cdot\,\|_{[R]}$ becomes, in some sense, closer to being translation and modulation invariant. From suitable limit process, letting $R$ tending to infinity, we are able to extract a sought equivalent Hilbert norm which is norm preserved under translations and modulations. (See Lemma 1.3.)
@@
The proofs of Theorem 0.3 and its extensions, are, among others, based on the fact that the involved Hilbert spaces are separable. In Section 2 we verify such facts, by using the Bargmann transform to transfer the questions to Hilbert spaces of entire functions. In the end we conclude that any Hilbert space which are continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$ (or in some even larger distribution spaces), must be separable. (See Propositions 1.2 and 2.2.)
@@ Acknowledgement
The first and second authors are grateful to HRI and the infosys foundations for support for the second author during his visit at HRI in December 2023, making this collaboration possible. The second author is also grateful for the excellent hospitality during his visit at HRI in December 2023. The second author was supported by Vetenskapsrådet (Swedish Science Council) within the project 2019-04890. The third author was supported by the Research Foundation–Flanders through the FWO-grant number G067621N.
@@ 1. Proof of the main result
In this section we prove our main result Theorem 0.3. First we show how the condition (0.2) in [2, Theorem 3.1] can be removed, which leads to Theorem 0.2. Then we prove a lemma, which in combination with Theorem 0.2, essentially leads to Theorem 0.3.
@@
Suppose that $\mathcal{H}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$. Then there is a constant $C>0$ such that (0.2) holds.
Lemma 1.1 follows by a straight-forward continuity argument. (See e. g. the arguments in [14, p.35].) In order to assist the reader we present a proof in Section 2 for a more general result (see Lemma 2.1 and its proof).
Proof of Theorem 0.2.
The result follows by combining [2, Theorem 3.1] with Lemma 1.1. The details are left for the reader. ∎
The proof of Theorem 0.3 is based on the fact that the involved Hilbert spaces are separable. This fact is guaranteed by the following proposition.
Proposition 1.2.
Suppose that the Hilbert space $\mathcal{H}$ is continuously embedded in $\mathscr{S}^{\prime}(\mathbf{R}^{d})$. Then $\mathcal{H}$ is separable.
We observe that neither translation nor modulation invariant hypotheses are imposed on the Hilbert spaces in Proposition 1.2.
It is expected that Proposition 1.2 is available in the literature. For completeness we give a proof of a generalized result in Section 2 (see Proposition 2.2 and its proof).
@@
Furthermore, $\|\,\cdot\,\|_{[R]}$ is a norm which arises from the scalar product
@@
exists and is independent of the chosen particular sequence $\{f_{0,j}\}_{j=1}^{\infty}$. Since (1.2) holds for any $f_{0,j}\in\mho$, it follows from the recent estimates and limit properties that (1.2) extends to any $f\in\mathcal{H}$. This gives (1).
@@
$\Delta_{R}=\big{(}(-x_{j}e_{j}+Q_{R})\setminus Q_{R}\big{)}\bigcup\big{(}Q_{R}\setminus(-x_{j}e_{j}+Q_{R})\big{)},$
@@
2. Separability of Hilbert spaces embedded in distribution spaces
In this section we give a proof of that Hilbert spaces which are continuously embedded in suitable distribution spaces are separable. We remark that these distribution spaces can be significantly larger than the set of tempered distributions, $\mathscr{S}^{\prime}(\mathbf{R}^{d})$. Especially we deduce a generalization of Proposition 1.2 (see Proposition 2.2 below).
First we introduce some test function spaces and their distribution spaces, which are under consideration. We recall that the Pilipović space $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ consists of all Hermite function expansions
@@ 2. Separability of Hilbert spaces embedded in distribution spaces
for some $h>0$. As in [16], we equip $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ with the inductive limit topology of $\mathcal{H}_{\flat,h}(\mathbf{R}^{d})$ with respect to $h>0$. Here $\mathcal{H}_{\flat,h}(\mathbf{R}^{d})$ is the Banach space of all smooth functions $f$ on $\mathbf{R}^{d}$ such that
$\|f\|_{\mathcal{H}_{\flat,h}}\equiv\sup_{\alpha\in\mathbf{N}^{d}}\left(\frac{|c(f,\alpha)|\alpha!^{\frac{1}{2}}}{h^{|\alpha|}}\right)$
is finite.
The distribution space $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ can be identified with the set of all formal expansions in (2.1) such that
@@
for every $h>0$. The topology of $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ is the projective limit topology of $\mathcal{H}_{\flat,h}^{\prime}(\mathbf{R}^{d})$ with respect to $h>0$. Here $\mathcal{H}_{\flat,h}^{\prime}(\mathbf{R}^{d})$ is the Banach space of all smooth functions $f$ on $\mathbf{R}^{d}$ such that
$\|f\|_{\mathcal{H}_{\flat,h}^{\prime}}\equiv\sup_{\alpha\in\mathbf{N}^{d}}\left(\frac{|c(f,\alpha)|\alpha!^{-\frac{1}{2}}}{h^{|\alpha|}}\right)$
is finite. It follows that $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ is complete and that $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ is a Fréchet space.
The distribution action between $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ and $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ is then given by
@@
with dense and continuous embeddings. Since $\overline{h_{\alpha}}=(-1)^{|\alpha|}h_{\alpha}$, the duality between $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ and $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ in (2.2) can also be expressed by
$\langle f,\phi\rangle=\sum_{\alpha\in\mathbf{N}^{d}}(-1)^{|\alpha|}c(f,\alpha)c(\phi,\alpha),$ (2.4)
and it follows that $\langle f,\phi\rangle$ agrees with the usual distribution actions between elements in $\mathscr{S}(\mathbf{R}^{d})$ and $\mathscr{S}^{\prime}(\mathbf{R}^{d})$, when $\phi\in\mathcal{H}_{\flat}(\mathbf{R}^{d})$ and, more restrictive, $f\in\mathscr{S}^{\prime}(\mathbf{R}^{d})$.
We are now prepared to state the generalizations of Lemma 1.1 and Proposition 1.2 in Section 1.
Lemma 2.1.
Suppose that $\mathcal{H}$ is continuously embedded in $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$. Then there is a constant $C>0$ such that (0.2) holds.
Proof.
Note that $p_{\phi}(f)=|(f,\phi)|$, $f\in\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$, is a continuous seminorm on $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$. The continuity of the inclusion mapping $\mathcal{H}\to\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$ then yields $p_{\phi}(f)\leq C_{\phi}\|f\|_{\mathcal{H}}$ for some $C_{\phi}>0$ and all $f\in\mathcal{H}$, which is the same as (0.2), completing the proof. ∎
We observe that the Hilbert structure of $\mathcal{H}$ actually plays no role in the previous proof and Lemma 2.1 therefore holds if we just assume that $\mathcal{H}$ is a Banach space that is continuously embedded in $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$.
Our generalization of Proposition 1.2 is the following.
Proposition 2.2.
Suppose that the Hilbert space $\mathcal{H}$ is continuously embedded in $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$. Then $\mathcal{H}$ is separable.
Evidently, by (2.3) it follows that Proposition 2.2 is true with $\Sigma_{1}^{\prime}$ in place of $\mathcal{H}_{\flat}^{\prime}$.
We need some preparations for the proof of Proposition 2.2. Especially we shall make use of the Bargmann transform, $\mathfrak{V}_{d}$, defined by
[…]
We have
[…]
or
$(\mathfrak{V}_{d}f)(z)=\langle f,\mathfrak{A}_{d}(z,\,\cdot\,)\rangle,\quad z\in\mathbf{C}^{d},$ (2.5)
where the Bargmann kernel $\mathfrak{A}_{d}$ is given by
$\mathfrak{A}_{d}(z,y)=\pi^{-\frac{d}{4}}\exp\Big{(}-\frac{1}{2}(\langle z,z\rangle+|y|^{2})+2^{1/2}\langle z,y\rangle\Big{)},\quad z\in\mathbf{C}^{d},y\in\mathbf{R}^{d}.$
(Cf. [3, 4].) Here
$\langle z,w\rangle=\sum_{j=1}^{d}z_{j}w_{j}\quad\text{and}\quad(z,w)=\langle z,\overline{w}\rangle$
when
$z=(z_{1},\dots,z_{d})\in\mathbf{C}^{d}\quad\text{and}\quad w=(w_{1},\dots,w_{d})\in\mathbf{C}^{d}.$
[…]
@@
are homeomorphisms.
Proof of Proposition 2.2.
We consider the Hilbert space of entire functions ~$\tilde{\mathcal{H}}=\mathfrak{V}_{d}(\mathcal{H})$, provided with the scalar product $(\mathfrak{V}_{d}f,\mathfrak{V}_{d}g)_{\tilde{\mathcal{H}}}=(f,g)_{\mathcal{H}}$. It thus suffices to show that $\tilde{\mathcal{H}}$ is separable. Let $z\in\mathbf{C}^{d}$. Reasoning as in the proof of Lemma 1.1 with the test function $\mathfrak{A}_{d}(z,\>\cdot\>)$ instead of $\phi$, we obtain the existence of a constant $C_{z}>0$ such that
$|\mathfrak{V}_{d}f(z)|=|\left\langle f,\mathfrak{A}_{d}(z,\>\cdot\>)\right\rangle|\leq C_{z}\|f\|_{\mathcal{H}}=C_{z}\|\mathfrak{V}_{d}f\|_{\tilde{\mathcal{H}}},\qquad f\in\mathcal{H},$
which shows that $\tilde{\mathcal{H}}$ is a reproducing kernel Hilbert space. (See e. g. [1].)
Let $K_{z}(w)=K(z,w)$ be its reproducing kernel and fix a dense subset $D$ of $\mathbf{C}^{d}$. We claim that the linear span of $\{\,K_{z}\,;\,z\in D\,\}$ is a dense subset of $\mathcal{H}$. Indeed, if $F\in\tilde{\mathcal{H}}$ is such that $F(z)=(F,K_{z})_{\tilde{\mathcal{H}}}=0$ for each $z\in D$, then $F$ must identically vanish in view of the density of $D$ and the continuity of $F$. ∎
Remark 2.3.
Let $d\mu(z)=e^{-|z|^{2}}d\lambda(z)$, $z\in\mathbf{C}^{d}$, where $d\lambda(z)$ is the Lebesgue measure on $\mathbf{C}^{d}$. Then the Bargmann-Fock space, $A^{2}(\mathbf{C}^{d})$, consists of all $F\in A(\mathbf{C}^{d})$ such that
[…]
is finite. We recall that $A^{2}(\mathbf{C}^{d})$ is a Hilbert space with scalar product
[…]
In [3] it is proved that if $\mathcal{H}$ and its norm, are equal to $L^{2}(\mathbf{R}^{d})$ and its norm, then $\tilde{\mathcal{H}}$ and its norm, are equal to $A^{2}(\mathbf{C}^{d})$ and its norm.
Let
$\mathcal{A}_{\flat}(\mathbf{C}^{d})\equiv\bigcup_{r\geq 0}\{\,F\in A(\mathbf{C}^{d})\,;\,|F(z)|\leq Ce^{r|z|}\ \text{for some}\ C>0\,\},$
equipped with the inductive limit topology through the semi-norms
$\|F\|_{\mathcal{A}_{\flat}\,,r}\equiv\|F\cdot e^{-r|\,\cdot\,|}\|_{L^{\infty}}.$
Evidently, $\mathcal{A}_{\flat}(\mathbf{C}^{d})$ is continuously embedded in $A^{2}(\mathbf{C}^{d})$. In [16] it is proved that
  1. (1)
    the Bargmann transform is a homeomorphism from $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ to $\mathcal{A}_{\flat}(\mathbf{C}^{d})$;
  2. (2)
    $\mathcal{A}_{\flat}(\mathbf{C}^{d})\subseteq A^{2}(\mathbf{C}^{d})\subseteq\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})$ with dense embeddings;
  3. (3)
    the map $(F,G)\mapsto(F,G)_{A^{2}}$ from $\mathcal{A}_{\flat}(\mathbf{C}^{d})\times\mathcal{A}_{\flat}(\mathbf{C}^{d})$ to $\mathbf{C}$ extends uniquely to a continuous map from $\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})\times\mathcal{A}_{\flat}(\mathbf{C}^{d})$ or from $\mathcal{A}_{\flat}(\mathbf{C}^{d})\times\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})$ to $\mathbf{C}$;
  4. (4)
    the (strong) dual of $\mathcal{A}_{\flat}(\mathbf{C}^{d})$ can be identified with $\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})$, through the form $(\,\cdot\,,\,\cdot\,)_{A^{2}}$. Furthermore,
    $(\mathfrak{V}_{d}f,\mathfrak{V}_{d}g)_{A^{2}}=(f,g)_{L^{2}},\qquad f\in\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d}),\ g\in\mathcal{H}_{\flat}(\mathbf{R}^{d}).$
The reproducing kernel of $A^{2}(\mathbf{C}^{d})$ is given by $K_{z}(w)=e^{(w,z)}$ (see e. g. [3]). Hence
$F(z)=(F,e^{(\,\cdot\,,z)})_{A^{2}},$ (2.9)
when $F\in A^{2}(\mathbf{C}^{d})$. We observe that $K_{z}\in\mathcal{A}_{\flat}(\mathbf{C}^{d})$, for every $z\in\mathbf{C}^{d}$. Hence, the right-hand side of (2.9) makes sense for any $F\in\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})$, and because $A^{2}(\mathbf{C}^{d})$ is dense in $F\in\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})$, it follows that the identity (2.9) still holds for any $F\in\mathcal{A}_{\flat}^{\prime}(\mathbf{C}^{d})$.
We observe that the reproducing kernel in the proof of Proposition 2.2 is chosen with respect to the scalar product $(\,\cdot\,,\,\cdot\,)_{\tilde{\mathcal{H}}}$, while the reproducing kernel in (2.9) is defined with respect to $(\,\cdot\,,\,\cdot\,)_{A^{2}}$. It follows that these kernels are not the same, when $\mathcal{H}$ differs from $L^{2}(\mathbf{R}^{d})$.
3. Extensions and variations
We have now the following extension of Theorem 0.3.
Theorem 3.1.
The proof of Theorem 3.1 follows by a combination of [18, Proposition 1.5], Theorem 0.3, and Proposition 3.3 below.
We need some preparations for the proof of Theorem 3.1. Since translation and modulation invariant Hilbert spaces are in focus, we here notice that all the spaces in Theorem 3.1 are invariant under such actions, which is explained in the next result.
Proposition 3.2.
Let $x,\xi\in\mathbf{R}^{d}$. Then the map $f\mapsto f(\,\cdot\,-x)e^{i\langle\,\cdot\,,\xi\rangle}$ is a homeomorphism on each one of the spaces in (2.3) and in $\mathscr{D}^{\prime}(\mathbf{R}^{d})$.
Proof.
The result is evidently true for the spaces
$\mathscr{S}(\mathbf{R}^{d}),\quad\Sigma_{1}(\mathbf{R}^{d}),\quad\mathscr{S}^{\prime}(\mathbf{R}^{d}),\quad\Sigma_{1}^{\prime}(\mathbf{R}^{d})\quad\text{and}\quad\mathscr{D}^{\prime}(\mathbf{R}^{d}).$
We need to prove the result for $\mathcal{H}_{\flat}(\mathbf{R}^{d})$ and $\mathcal{H}_{\flat}^{\prime}(\mathbf{R}^{d})$.
We recall that in terms of the Weyl map